# Determinations of angular stiffness in rotational optical tweezers

### v1

@@ III Determination of stiffness from angular trajectories
where $I=(2/5)ma^{2}$ is the inertia of the probe (mass $m$, radius $a$), $\gamma_{0}=8\pi\eta a^{3}$ is the Stokes rotational drag coefficient for a sphere in a viscous medium, $k_{B}$ is Boltzmann’s constant, $T$ is the temperature, and $\xi(t)$ is a Gaussian white-noise term with with $\langle\xi(t)\rangle=0$, $\langle\xi(t)^{2}\rangle=1$, and $\langle\xi(t)\xi(t^{\prime})\rangle=\delta(t-t^{\prime})$.
The right hand side of Eq. 4 represents the rotational Brownian motion that drives angular fluctuations about equilibrium. The rate at which the angular trajectory returns to the equilibrium is set by the characteristic trap time, $\tau_{\mathrm{opt}}=\gamma_{0}/\chi$, which is the ratio of the drag coefficient to the angular trap stiffness. Hence, the calibration techniques often analyse the temporal or frequency correlations of the angular trajectory to determine $\tau_{\mathrm{opt}}$ as an intermediate step before obtaining $\chi$.
Typically, the inertia dissipation occurs significantly faster than the experimental measurement rate (on the order of ${10}^{-6}\text{\,}\mathrm{s}$), such that $I\ddot{\phi}(t)\simeq 0$. The angular trajectory can be approximated as
@@
To extract the angular trap stiffness from experimental trajectories, we employ several complementary analysis techniques that exploit different statistical properties of the angular fluctuations and are frequently used in the calibration of translational trap stiffness. These include equipartition (EQP), mean‑squared displacement (MSD), autocorrelation function (ACF), power spectral density (PSD), and maximum likelihood estimation (MLE) methods. Each provides a complementary route to determine $\chi$, with distinct advantages and sensitivities to experimental artefacts such as finite sampling, calibration accuracy, or detector noise. By applying all five approaches to the same dataset, we both validate the consistency of the extracted stiffness and highlight the practical considerations that arise in their implementation.
@@ III.1 Equipartition (EQP) Analysis
The simplest approach to determine the angular trap stiffness is through the equipartition theorem, which relates the thermal energy of the system to the variance of the angular fluctuations on timescales much larger than the characteristic trapping time. In the harmonic approximation, this becomes
$\frac{1}{2}\chi\langle\phi(t)^{2}\rangle=\frac{1}{2}k_{B}T.$ (7)
A direct expression of $\chi$ can be obtained as
$\chi_{\mathrm{\ignorespaces\penalty 10000\ }}=\frac{k_{B}T}{\langle\phi(t)^{2}\rangle}.$ (8)
In practice, $\langle\phi(t)^{2}\rangle$ is obtained from the experimentally measured angular trajectory after subtracting any slow drift or offset. This method is straightforward and requires only a measurement of angular position, with no knowledge of the local viscosity, making it attractive for rapid estimation of $\chi$. Its accuracy critically depends on the precision of the angular calibration, and the validity of the approach relies on the assumption that the probe remains in the linear regime of the trap, and that a sufficiently long measurement time to adequately sample the distribution has been made. The histograms in Fig. 2A show the count of angular position for a probe at three different trapping powers. Higher trapping powers lead to stiffer traps, resulting in a smaller variance. Fig. 2B presented the same data on a logarithmic scale, fitted with a quadratic function to obtain $\chi$. Only the central 99% of the data was used, as the counts at larger angular displacements were insufficient for a reliable average.
@@ III.2 Mean-squared displacement (MSD) Analysis
The mean-squared displacement of the angular trajectory offers an alternative approach to determine $\chi$. The MSD of a particle, $\langle\left(\phi(t+\tau)-\phi(\tau)\right)^{2}\rangle$, quantifies the average displacement a particle can travel given some lag time, $\tau$. A free particle undergoing rotational diffusion has an MSD $\propto\tau$. The angular confinement from the optical trap prevents the particle from freely diffusing and introduces an exponential decay to the MSD, which plateaus at time scales longer than $\tau_{\mathrm{opt}}$.
The theoretical MSD can be derived using Eq. 6 as follows:
@@
$=\left\langle\phi(t+\tau)^{2}\right\rangle-2\left\langle\phi(t+\tau)\phi(t)\right\rangle+\left\langle\phi(\tau)^{2}\right\rangle,$ (10)
@@
where the integral for $s\in[t,t+\tau]$ vanishes because the noise term, $\xi(s)$, is independent of $\phi(t)$ in this domain and has a mean of zero. Therefore, the expected MSD becomes
@@
As expected, $\mathrm{MSD}_{\tau\ll\tau_{\text{opt}}}\propto\tau$ corresponding to free rotational diffusing, while $\text{MSD}_{\tau\gg\tau_{\text{opt}}}$ is the the maximum displacement reached by the particle.
Fig. 2C shows the experimental MSD calculated at specific lag times obtained for three different trapping powers, together with fits to the theoretical form (Eq. 14). Increasing trap power leads to a faster saturation and a lower plateau, consistent with a stiffer angular potential. Fitting the experimental MSD data to the theoretical expression yields both $\tau_{\mathrm{opt}}$ and $\chi$, allowing a simultaneous estimation of $\gamma_{0}$ and stiffness.
This method is robust to slow drift and does not require explicit calibration of the torque signal. However, it is sensitive to finite trajectory length and sampling rate, which can bias the long‑lag plateau. When calculating the MSD, either overlapping (evaluating the displacement for every possible time origin) or non‑overlapping intervals can be used. The overlapping approach provides a much larger number of data points for averaging, allowing the MSD curve to be resolved with shorter measurement times. However, successive MSD values are not strictly independent introducing correlations between points. By contrast, non‑overlapping calculations avoid this correlation but require substantially longer trajectories to achieve the same statistical precision.
@@ III.3 Autocorrelation function (ACF) Analysis
The autocorrelation function (ACF) provides a complementary description of the angular dynamics that is linked to the MSD. While the MSD emphasises the growth of angular fluctuations over some lag time, the ACF highlights the decay of correlations back to equilibrium. Angular confinement ensures that positions separated by short time intervals remain correlated. The correlation is given by:
@@
where the solution follows directly from the corresponding term in the MSD derivation (Eq. 13). In fact, the MSD and ACF quantities are connected through $\mathrm{MSD}(\tau)=2\left(\langle\phi^{2}\rangle-C_{\phi}(\tau)\right)$. For a discrete trajectory, the correlation is estimated as
@@
Fig. 2D shows the measured ACFs for three different trapping powers, together with exponential fits to the theoretical form. Increasing trap power leads to a faster decay of correlations, consistent with a stiffer angular potential. This method is less sensitive to slow drift than equipartition and can resolve the relaxation time from shorter trajectories than the MSD. Although, it can be affected by finite sampling, similar to the MSD method.
@@ III.4 Power Spectral Density (PSD) Analysis
The power spectral density (PSD) provides a frequency‑domain description of the thermally driven angular fluctuations. PSD analysis has become one of the most common methods in optical trapping and microrheology, where it is routinely used to extract trap stiffness and hydrodynamic parameters.
It is mathematically related to the autocorrelation function through the Wiener–Khinchin theorem: the PSD is the Fourier transform of the ACF [d_c_champeney_handbook_1987]. The PSD emphasises how the variance of the signal is distributed across frequencies, while the ACF highlights the temporal decay of those correlations. Both approaches contain equivalent information, but the PSD is often preferred in practice due to its robustness against drift and its ability to reveal noise characteristics across a wide frequency range. Computationally, the fast Fourier transform (FFT) algorithms make obtaining the PSD inexpensive compared with directly calculating the autocorrelation for an equivalent number of lag times. In many cases, it is more efficient to compute the ACF by taking the inverse Fourier transform of the PSD.
The power spectrum models the expected behaviour of the PSD. . The angular position power spectrum (APPS) is defined as $S_{\phi}(f)=|\tilde{\phi}(f)|^{2}$, where the tilde represents the unilateral Fourier transform. The Fourier transform of Eq. 5,
@@
where $f_{c}=\chi/(2\pi\gamma_{0})=1/(2\pi\tau_{\mathrm{opt}})$ is the corner frequency, with white noise of power, $|\tilde{\xi}(f)|^{2}=1$. The power spectrum takes the form of a Lorentzian, with a flat plateau at low frequencies consistent with harmonic motion in a high loss medium and a $f^{-2}$ proportionality of decay at high frequencies that is characteristic of Brownian walks [berg-sorensen_power_2004]. The corner frequency is the inverse of the characteristic angular trapping time constant of the probe in the fluid, and it marks the transition between these regimes, providing a direct measure of the trap stiffness. In an experiment, we cannot find the power spectrum, but instead find the PSD, which is obtained with $S_{\phi}(f)=|\tilde{\phi}(f)|^{2}/t_{\mathrm{meas}}$, where $t_{\mathrm{meas}}$ is the total measurement time.
Fitting the PSD to Eq. 20 often yields both $D$ and $f_{c}$ from which a robust estimate of $\gamma_{0}$ and $\chi$ are obtained. The experimental noise spectrum can be measured and subtracted from the PSD before fitting to improve accuracy. Fig. 2E shows the PSDs for three different trapping powers, together with Lorentzian fits. Increasing trap power shifts the corner frequency to higher values, consistent with a stiffer angular potential. The data sets are truncated at the Nyquist frequency to minimise aliasing. Windowing and segment averaging are commonly used to reduce spectral leakage and improve statistical reliability. Deviations from the expected dynamics due to the hydrodynamic and inertial effects are discussed later in Section V.3 and can be neglected for most rotational experiments outside the rotational ballistic regime [watson_interrogating_2025].
@@ III.5 Maximum Likelihood Estimation (MLE)
Maximum likelihood estimation is an alternative approach to determining the angular trap stiffness. MLE has been employed in translational measurements—by analysing the position or force at sequential time points—to calibrate the translational trap stiffness, diffusion constant, and for 3D reconstruction for the optical potential[perez_garcia_high-performance_2018, stilgoe_enhanced_2021, Gieseler2021]. In contrast, MLE estimation for angular trap stiffness determination can be achieved without any temporal information by combining two separate measurements of optical torque.
@@
that can be maximised with respect to $\chi$. Taking $\partialderivative{L}{\chi}=0$, gives a closed form-estimator
@@
for the angular trap stiffness. Interestingly, this approach can be modified to estimate the conversion factor from detector $D_{0}$ in Fig. 1 to the optical trapping power if $\chi$ is already known [watson_rotational_2022].
@@
$\Delta\phi_{i}-\omega_{c}\phi_{i}=\sqrt{\frac{2D}{\Delta t}}\xi_{i},$ (24)
where $\omega_{c}=\chi/\gamma=1/\tau_{\mathrm{opt}}$ is the angular corner frequency. Following previous derivations in the translational regime [perez_garcia_high-performance_2018, stilgoe_enhanced_2021], $\omega_{c}$ is estimated using MLE, by
$\omega_{c}=\frac{\sum_{i}^{N}(\phi_{i+1}-\phi_{i})(\phi_{i})}{\sum_{i}^{N}\phi_{i}^{2}}.$ (25)
@@
$D=\frac{\Delta t}{2N}\sum_{i}^{N}(\phi_{i+1}+(\omega_{c}-1)\phi_{i}).$ (26)
@@ III.6 Summary of passive calibration methods
So far, we see that the calibration results from rotational optical tweezers experiments shadow those of regular optical tweezers. However, these similarities end at the assumption of a linear model. The mode of measurement and the differences in how angular momentum is transferred become more important in the following sections.
@@ IV Numerical Determination of Stiffness
Figure 3: (A) The force profile for a $1\lambda$ radius probes. (B) The torque profile for a $1\lambda$ radius vaterite showing rotation about each axis. (C) The torque profile about $z$ for three differently sized probes. (D) The calculated $\chi$ for a given radius using the torque profile method. (E) Heatmap of the angular trap stiffness relative to beam size and beam convergence angle for a vaterite probe. (F) A heatmap of the optical torque on a vaterite probe when illuminated in circularly polarised light. In both (E) and (F), only the conditions where a stable spatial trap occurs are plotted, with the probe positioned at its equilibrium position along the beam axis as calculated from the force profile.
Having established consistent estimates of the angular trap stiffness from multiple experimental approaches, we will use numerical simulations to investigate how the angular trap stiffness varies from calibrations away from the linear stiffness limit. Computational methods offer two complementary routes to achieve this. Firstly, angular trajectories can be simulated free from instrumental drift, aliasing, and detector noise, allowing us to isolate the intrinsic stochastic dynamics of the trapped probe. These results can serve as a benchmark for the stochastic estimates. Secondly, the angular trap stiffness can be computed directly by calculating the transfer of angular momentum to the probe particle as its angular displacement $\phi$ is varied. This methodology enables systematic exploration of different trapping configurations and the influence of specific parameters on the angular trap stiffness in controlled systems.
The following numerical results were obtained using the T‑matrix method implemented in the Optical Tweezers Toolbox [OTT2007] to calculate the optical force and torque on the probe particle [Loke2009, Nieminen2011]. The translational trap stiffness and equilibrium position can be obtained from the force profile, i.e. the force-displacement curve along each axis. Fig. 3A shows the force profiles for a probe of radius $1\lambda$. Owing to to the rotational symmetry of the beam, spherical particles often have their equilibrium position ($x_{0},y_{0},z_{0}$) on the beam axis. Consequently, the force profile is first calculated along the z-axis to determine the axial trapping location. The local slope of the force–displacement curve about this equilibrium in the $\hat{x}$, $\hat{y}$, and $\hat{z}$ directions then defines the corresponding trap stiffnesses, $\kappa_{x}$, $\kappa_{y}$, and $\kappa_{z}$, respectively.
A similar approach can be applied to each rotational degree of freedom, with the torque profiles shown in Fig. 3B. There is an alignment torque in the azimuthal and out-of-plane rotations, which can be used to define $\chi_{\phi}$ and $\chi_{\theta}$, respectively. Unlike the azimuthal dynamics, there is an alignment torque that keeps the probe’s optic axis orthogonal to the beam propagation direction. As Fig. 3B demonstrates, there is no alignment torque about the probe’s optic axis due to the rotational symmetry of its anisotropy, hence $\chi_{\rho}=0$.
As with the translational trap stiffness, the angular trap stiffness can be obtained from the local slope of the torque profile at equilibrium. The azimuthal torque profiles are shown in Fig. 3C for three probe sizes, with a solid line indicating the local slope used to extract $\chi_{\phi}$. The dependence of the angular trap stiffness on probe size is summarised in Fig. 3D. As expected from the scaling of optical torque with particle cross‑section, the angular stiffness increases with radius in a given trap.
@@
The numerical results so far have been calculated for a beam focussed by a NA 1.2 water-immersion objective, corresponding to the objective lens in our ROTs system. However, computational results can be used to investigate the angular trap stiffness with a variety of experimental parameters. Fig. 3E maps the influence of probe size and beam convergence angle on the angular trap stiffness. As the convergence angle increases, there is a greater transfer of spin angular momentum due to the larger scattering cross-sections and the shift in the axial equilibrium position toward the beam focus. The optical torque is maximised at the beam focus, so the angular stiffness is sensitive to changes in the axial equilibrium. At low convergence angles a stable trap may not form, which appears as empty regions in the heat map. The boundary of this unstable region oscillates with probe radius due to interference effects [Stilgoe2008].
@@
For comparison, Fig. 3F presents a heat map of the optical torque dependent on probe size and beam convergence angle for a probe in a circularly polarised beam. In this case, the optical torque about the beam axis is independent of $\phi$, but its magnitude varies systematically with the convergence angle and the axial equilibrium position. The overall trend mirrors that observed for the angular stiffness in linearly polarised traps. Notably, the boundaries of the stable trapping region shift slightly under circular polarisation, reflecting the altered balance between gradient and scattering forces compared with the linearly polarised case.

### v2

@@ III Passive Methods for Angular Stiffness Determination
where $I=(2/5)ma^{2}$ is the inertia of the probe (mass $m$, radius $a$), $\gamma_{0}=8\pi\eta a^{3}$ is the Stokes rotational drag coefficient for a sphere rotating in a viscous medium, $k_{B}$ is Boltzmann’s constant, $T$ is the temperature, and $\xi(t)$ is a Gaussian white-noise term with with $\langle\xi(t)\rangle=0$, $\langle\xi(t)^{2}\rangle=1$, and $\langle\xi(t)\xi(t^{\prime})\rangle=\delta(t-t^{\prime})$.
The right hand side of Eq. 4 represents the rotational Brownian motion that drives angular fluctuations about equilibrium. Hence, calibration techniques typically require analysing the temporal or spatial correlations within the angular trajectory. The rate at which the angular displacements return to the equilibrium is given by the characteristic trap time, $\tau_{\mathrm{opt}}=\gamma_{0}/\chi$, which is the ratio of the drag coefficient to the trap stiffness.
The inertia dissipation occurs significantly faster than most experimental measurement rates (on the order of ${10}^{-6}\text{\,}\mathrm{s}$), such that $I\ddot{\phi}(t)\simeq 0$. Here, the angular trajectory becomes
@@
To extract the angular trap stiffness from experimental trajectories, we employ several analysis techniques that are frequently used in the calibration of translational trap stiffness. These include equipartition (EQP), mean‑squared displacement (MSD), autocorrelation function (ACF), power spectral density (PSD), and maximum likelihood estimation (MLE) methods. Each provides a complementary route to determine $\chi$, with distinct advantages and sensitivities to experimental artefacts such as finite sampling, calibration accuracy, or detector noise. These methods are valid provided the angular trajectory remain in the linear regime and there is a sufficiently long measurement time to resolve $\tau_{\mathrm{opt}}$.
@@ III.1 Equipartition (EQP) analysis
The simplest approach to determine the angular trap stiffness is through the equipartition theorem, which relates the thermal energy of the system to the variance of the angular fluctuations on timescales much larger than the characteristic trapping time. The linear restoring torque produces a harmonic potential, such that a direct expression of $\chi$ can be obtained as
$\chi=\frac{k_{B}T}{\langle\phi(t)^{2}\rangle}.$ (7)
In practice, $\langle\phi(t)^{2}\rangle$ is obtained after subtracting any slow drift or offset from the angular trajectory. This method only requires a measurement of angular position, with no knowledge of the local viscosity, making it attractive for rapid estimation of $\chi$. The method requires a sufficiently long measurement time to adequately sample the distribution has been made. The histograms in Fig. 2A show the count of angular position for a probe at three different trapping powers. Higher trapping powers lead to stiffer traps, resulting in a smaller variance. Fig. 2B presented the same data on a logarithmic scale, fitted with a quadratic function to obtain $\chi$. Only the central 99% of the data was used, as the counts at larger angular displacements were insufficient for a reliable average.
@@ III.2 Mean-squared displacement (MSD) analysis
The MSD of a particle quantifies the average displacement a particle can travel given some lag time, $\tau$.The MSD of a free particle undergoing rotational diffusion scales linearly with $\tau$. The angular confinement from the optical trap prevents the particle from freely diffusing and introduces an exponential decay to the MSD, which plateaus at time scales significantly longer than $\tau_{\mathrm{opt}}$.
The MSD is defined by
@@
$=\left\langle\phi(t+\tau)^{2}\right\rangle-2\left\langle\phi(t+\tau)\phi(t)\right\rangle+\left\langle\phi(\tau)^{2}\right\rangle.$ (9)
@@
The integral for $s\in[t,t+\tau]$ vanishes because the noise term, $\xi(s)$, is independent of $\phi(t)$ in this domain and has a mean of zero. Therefore, the expected MSD becomes
@@
As expected, $\mathrm{MSD}_{\tau\ll\tau_{\text{opt}}}\propto\tau$ corresponding to free rotational diffusion at short time scales, while $\text{MSD}_{\tau\gg\tau_{\text{opt}}}$ describes the plateau of the MSD at long times.
Fig. 2C shows the experimental MSD obtained for three different trapping powers, together with fits to the expected form in Eq. 13. Increasing trap power leads to a faster saturation and a lower plateau, consistent with a stiffer angular potential. Fitting the experimental MSD data to the theoretical expression yields both $\tau_{\mathrm{opt}}$ and $\chi$, allowing a simultaneous estimation of $\gamma_{0}$ and stiffness. This method is robust to slow drift and does not require explicit calibration of the torque signal. However, it is sensitive to finite trajectory length and sampling rate, which can bias the long‑lag plateau.
When calculating the MSD, either overlapping (evaluating the displacement for every possible time origin) or non‑overlapping intervals can be used. The overlapping approach provides a much larger number of data points for averaging, allowing the MSD curve to be resolved with shorter measurement times. However, successive MSD values are not strictly independent, introducing correlations between points. By contrast, non‑overlapping calculations avoid this correlation but require substantially longer trajectories to achieve the same statistical precision.
@@ III.3 Autocorrelation function (ACF) analysis
The ACF provides a complementary description of the angular dynamics that is linked to the MSD. While the MSD emphasises the growth of angular fluctuations over lag time, the ACF highlights the decay of correlations. Angular confinement ensures that positions separated by short time intervals remain correlated. The correlation is given by:
@@
where the solution follows directly from the corresponding term in the MSD derivation (Eq. 12). The MSD and ACF quantities are explicitly connected through $\mathrm{MSD}(\tau)=2\left(\langle\phi^{2}\rangle-C_{\phi}(\tau)\right)$. For a discrete trajectory, the correlation is estimated as
@@
Fig. 2D shows the measured ACFs for three different trapping powers and fitted by the expected model in Eq. 16. Increasing trap power leads to a faster decay of correlations, consistent with a stiffer angular potential. This method is less sensitive to slow drift than equipartition and can resolve the relaxation time from shorter trajectories than the MSD. It can be affected by finite sampling similar to the MSD method.
@@ III.4 Power spectral density (PSD) analysis
The power spectral density (PSD) provides a frequency‑domain description of the thermally driven angular fluctuations. PSD analysis has become one of the most common methods in optical trapping and microrheology, where it is routinely used to extract trap stiffness and hydrodynamic parameters [4, 5, 15].
It is mathematically related to the autocorrelation function through the Wiener–Khinchin theorem: the PSD is the Fourier transform of the ACF [11]. The PSD emphasises how the variance of the signal is distributed across frequencies, while the ACF highlights the temporal decay of those correlations. Both approaches contain equivalent information, but the PSD is often preferred in practice due to its robustness against drift and its ability to reveal noise characteristics across a wide frequency range. Computationally, the fast Fourier transform (FFT) algorithms make obtaining the PSD inexpensive compared with directly calculating the autocorrelation for an equivalent number of lag times. In many cases, it is more efficient to compute the ACF by taking the inverse Fourier transform of the PSD.
The power spectrum models the expected behaviour of the PSD. The angular position power spectrum (APPS) is defined as $S_{\phi}(f)=|\tilde{\phi}(f)|^{2}$, where the tilde represents the unilateral Fourier transform. The Fourier transform of Eq. 5,
@@
where $f_{c}=\chi/(2\pi\gamma_{0})=1/(2\pi\tau_{\mathrm{opt}})$ is the corner frequency, with white noise of power, $|\tilde{\xi}(f)|^{2}=1$. The power spectrum takes the form of a Lorentzian, with a flat plateau at low frequencies consistent with harmonic motion in a high loss medium and a $f^{-2}$ proportionality of decay at high frequencies that is characteristic of Brownian walks [4]. The corner frequency is the inverse of the characteristic angular trapping time, marking the transition between these regimes and providing a direct measure of the trap stiffness. In an experiment, we cannot find the power spectrum, but instead find the PSD, which is obtained with $S_{\phi}(f)=|\tilde{\phi}(f)|^{2}/t_{\mathrm{meas}}$, where $t_{\mathrm{meas}}$ is the total measurement time.
Fitting the PSD to Eq. 19 often yields both $D$ and $f_{c}$ from which a robust estimate of $\gamma_{0}$ and $\chi$ are obtained. The experimental noise spectrum can be measured and subtracted from the PSD before fitting to improve accuracy. Fig. 2E shows the PSDs for three different trapping powers, together with Lorentzian fits. Increasing trap power shifts the corner frequency to higher values, consistent with a stiffer angular potential. The data sets are truncated at the Nyquist frequency to minimise aliasing. Windowing and segment averaging are commonly used to reduce spectral leakage and improve statistical reliability. Deviations from the expected dynamics due to the hydrodynamic and inertial effects are discussed later in Section V.3 and can be neglected for most rotational experiments outside the rotational ballistic regime [44].
@@ III.5 Maximum likelihood estimation (MLE)
Maximum likelihood estimation is an alternative approach to determining the angular trap stiffness. MLE has been employed in translational measurements—by analysing the position or force at sequential time points—to calibrate the translational trap stiffness, diffusion constant, and for 3D reconstruction for the optical potential [33, 37, 15]. In contrast, MLE estimation for angular trap stiffness determination can be achieved without any temporal information by combining two separate measurements of optical torque.
@@
that can be maximised with respect to $\chi$. Taking $\partialderivative{L}{\chi}=0$, gives a closed form-estimator of
@@
for the angular trap stiffness. Similar to the equipartition method, this formulation using the maximum likelihood estimation is independence of the probe size and the fluid properties, provided the expected linear dynamics hold true. Interestingly, this approach can be modified to estimate the conversion factor from detector $D_{0}$ in Fig. 1 to the optical trapping power if $\chi$ is already known [43].
@@
$\Delta\phi_{i}+\omega_{c}\phi_{i}=\sqrt{\frac{2D}{\Delta t}}\xi_{i},$ (23)
where $\omega_{c}=\chi/\gamma=1/\tau_{\mathrm{opt}}$ is the angular corner frequency. Following analogous derivations in the translational regime [33, 37], $\omega_{c}$ is estimated using MLE, by
$\omega_{c}=-\frac{1}{\Delta t}\frac{\sum_{i}^{N}(\phi_{i+1}-\phi_{i})(\phi_{i})}{\sum_{i}^{N}\phi_{i}^{2}}.$ (24)
@@
$D=\frac{1}{2N\Delta t}\sum_{i}^{N}(\phi_{i+1}+(\omega_{c}\Delta t-1)\phi_{i})^{2}.$ (25)
@@ III.6 Summary of passive calibration methods
[…]
@@ IV Numerical Determination of Stiffness
Figure 3: Numerical calculation of the (A) force and (B) torque profiles of a $1\lambda$ radius vaterite probein an $x$-polarised trapping beam (NA 1.2). (C) The torque profile about $z$ for three differently sized probes, with the linear approximation highlighted for angular displacements $15^{\circ}$. (D) The calculated $\chi$ for a given radius using the torque profile method. (E) Heatmap of the angular trap stiffness relative to beam size and beam convergence angle for a vaterite probe. (F) A heatmap of the optical torque on a vaterite probe when illuminated in circularly polarised light. In both (E) and (F), the probe is located at its equilibrium position. The empty regions of the heatmaps correspond to the conditions where no stable spatial trap is formed.
This section considers numerical methods to determining the angular trap stiffness and investigate its sensitivity to experimental parameters, such as probe size and the trapping beam’s properties. The computational approach allows isolation of the intrinsic stochastic dynamics without experimental noise sources or drift, exact control of experimental parameters, and direct calculations of optical forces and torques. The following results were obtained using the T‑matrix method implemented in the Optical Tweezers Toolbox [28], as described in section II.2.
The translational trap stiffness and equilibrium position can be obtained from the force profile, i.e. the force-displacement curve along each axis. These are displayed in Fig. 3A for a $1\lambda$ radius probe, with the local slope at the probe’s equilibrium position in the $\hat{x}$, $\hat{y}$, and $\hat{z}$ directions used to define the trap stiffnesses, $\kappa_{x}$, $\kappa_{y}$, and $\kappa_{z}$, respectively according to the linear restoring force model. The same approach is used to define the angular trap stiffnesses, $\chi_{x}$, $\chi_{y}$, and $\chi_{z}$, from the torque profiles (torque-azimuthal displacement curves) which are shown in Fig. 3B.
The linear model is not appropriate for calculating the translational stiffness in $\hat{x}$ or the angular stiffness about $\hat{x}$—more generally, in or about the polarisation axis. The isotropic internal structure of the vaterite causes the $x$ force profile to flatten out around the equilibrium. This is a size-dependent phenomena, that in some cases shifts the equilibrium position away from the beam axis [45], and requires a non-linear model that is not discussed in this work. Similarly, the rotational symmetry about the polarisation axis means the is no torque about $x$ and therefore $\chi_{x}=0$.
In the context of experiments, we are primarily interested in the rotational dynamics about the $z$-axis (azimuthal angle, $\phi$) as the polarisation detection scheme is sensitive to this mode of motion. The azimuthal torque profiles are shown in Fig. 3C for three probe sizes, with a solid line indicating the local slope used to extract $\chi$.
Fig. 3D demonstrates the dependence of the angular trap stiffness on probe size. The angular trap stiffness increases with probe size, as expected from the scaling of spin angular momentum transfer with particle cross‑section. A maximum angular trap stiffness of $1\penalty 10000\ \hbar/\mathrm{photon}/\mathrm{rad}$ is possible from the transfer of spin angular momentum and occurs where the probe effectively acts as a quarter waveplate. Vaterite microspheres reach this point near a radii of around $3\lambda$ [30].
@@
The numerical results so far have been calculated for a beam focussed by a NA 1.2 water-immersion objective, corresponding to the objective lens in our ROTs system. Fig. 3E maps the influence of probe size and beam convergence angle on the angular trap stiffness. As the convergence angle increases, there is a greater transfer of spin angular momentum due to the larger scattering cross-sections and the shift in the axial equilibrium position toward the beam focus. The optical torque is maximised at the beam focus, so the angular stiffness is sensitive to changes in the axial equilibrium. At low convergence angles a stable trap may not form, which appears as empty regions in the heat map. The boundary of this unstable region oscillates with probe radius due to interference effects [38].
@@
For comparison, Fig. 3F presents a heat map of the optical torque dependent on probe size and beam convergence angle for a probe in a circularly polarised beam. In this case, the optical torque about the beam axis is independent of $\phi$. The magnitude varies significantly with probe sizes and slightly with the convergence angle and the axial equilibrium position. The overall trend mirrors that observed for the angular stiffness in linearly polarised traps. Notably, the boundaries of the stable trapping region shift slightly under circular polarisation, reflecting the altered balance between gradient and scattering forces compared with the linearly polarised case.